# Primal-Dual iLQR for GPU-Accelerated Learning and Control in Legged Robots

### v1

@@ III Model Predictive Control
In this section, we describe briefly the dynamics model $\boldsymbol{h}(\boldsymbol{x}_{i},\boldsymbol{u}_{i})$ presented in (1b), introducing two different models used for the formulation of the OCP presented in the result Sec. IV. The efficacy of the control input obtained from a MPC strategy is directly related to the rate at which such a control strategy can be re-planned and to the accuracy of its predictions, as shown by [37]. Therefore, we consider two main formulations from the literature: one based on the Single Rigid Body Dynamics (SRBD) model and one based on the Whole Body (WB) model. The first approximates the system as a point mass with constant inertia, discarding the limb joint positions and the inertial variation from different limb configurations. This model is particularly effective in quadruped robots, where the limb mass is, in most cases, designed to be less than 10% of the total mass [38]. On the other hand, the latter completely captures the dynamics of the robot, enabling the maximum exploitation of the system’s capabilities. This benefit comes at the cost of handling a larger state space and highly non-convex models, thus affecting the maximum frequency at which the optimization problem can be solved. Moreover, the SRBD-MPC state is limited to the center of mass linear and angular positions and velocities. This means that, when representing the robot’s orientation using quaternions, the state lies in $\mathbb{R}^{7}$. For such model, the control input consists only of the Ground Reaction Forces (GRF), and thus lies in $\mathbb{R}^{3n_{c}}$, where $n_{c}$ is the number of contact points.
In contrast, the WB-MPC also includes joint positions and velocities in the state, and joint torques in the control input. This increases the dimensionality to $\mathbb{R}^{7+2n_{\text{joint}}}$ for the state and $\mathbb{R}^{3n_{c}+n_{\text{joint}}}$ for the control input.
@@ IV Result
Figure 2: Comparison of the MPC optimal cost, evaluated from the same initial condition with acados [8] (in blue) and with our implementation (in orange). Both solvers are evaluated at each call of the MPC and the costs are recorded.
(a)      Robots Trajectories
(b)      CoM Distance
Figure 3: a) The resulting path of the two robots tracking desired trajectories that will lead to a collision. b) the value of the distance between the CoMs of the two robots over time. The centralized controller tracks the trajectory while avoiding collision between the robots.
In this section, we present the results of various analyses that highlight the strength of the proposed approach. We evaluate the performance of our algorithm using the two models introduced in Sec. III (the SRBD and WB models). To benchmark our implementation against state-of-the-art solvers, we used acados [8] for comparisons on the SRBD model and crocoddyl [5] for the WB model.We used acados for the SRBD comparison, as it could be easily adapted for the evaluations with multiple robots 5 6. On the other hand, we used crocoddyl for the horizon length comparison with the WB model, as it showed significantly better performance than acados with the more complex model.
Our analysis includes investigating how the average solving time scales with the horizon length across all models and solvers. Additionally, we assess the solver’s performance as the system state dimension increases, as well as the vectorization capabilities of our approach. All tests were conducted on a desktop computer equipped with an Intel Core i7-13700KF and an NVIDIA RTX 3080. We used Mujoco XLA, to evaluate the WB dynamics. In all the presented benchmarks, if not stated differently, we used the 15 kg Unitree Go2 robot, a torque-controlled quadruped platform, for the ease of comparison with the state of the art. As Fig. 1 and the accompanying video  [^1] https://youtu.be/zquKLxbAU_Y show, our WB-MPC can also be implemented for real-time control of other robot morphologies, like humanoids.
(a)        SRBD Model
(b)        WB Model
Figure 4: Average solving times when varying the horizon length (number of nodes) for the SRBD model (top) and WB model (bottom). In blue, we plot the solution time of two different state-of-the-art solvers (acados [8] on the top, Crocoddyl [5] on the bottom), while in orange, we plot the one of our implementation.
@@ IV-A Performance evaluation
First, we evaluate the ability of our solver to generate complex dynamic maneuvers. In this scenario, we are generating a barrel roll motion with the Aliengo robot, as shown in Fig. 1. Thanks to its multiple shooting formulation, we can initialize the solver with an infeasible initial guess, and the solver converges in just $27$ iterations to a solution that completes the task. We also assess the performance of the proposed solver in terms of solution optimality. In Fig. 2, we compared the values of the cost from the optimal solution given by our implementation against the one given by acados, which is chosen for this comparison because it allows using the same cost function as the one in our formulation. In this scenario, we are using the whole-body model and are controlling the robot in simulation at 50Hz. The robot is tasked to trot with a forward speed of $0.3$ m/s while being randomly pushed by an external disturbance of $50$N for $0.25$s. At every control loop, we record the value of the cost functional from both solvers, comparing solutions that start from the same initial condition. Figure 2 reports the values recorded during a $2.5$s time span. Although the two solutions are comparably similar, our implementation results in a $20\%$ reduction in the average cost along the trajectory.
To finally validate the MPC, we also performed experiments on the real hardware, as shown in Fig. 1 and in the accompanying video. For the real experiments, the solver runs on an Nvidia RTX 4050 laptop GPU. Figure 1 shows snapshots of the robot walking blindly up the steps and through uneven terrain using only proprioceptive data for state estimation [39]. We also evaluated the proposed approach when used to control two different robots at the same time in a centralized control implementation in simulation. In this example, the two robots are controlled by the same MPC that evaluates the control input for both systems at $50$Hz. The MPC also includes a collision avoidance constraint, in the form of a quadratic penalty term. In the tested scenario, the two robots are tasked to track two perpendicular trajectories that cross each other and would lead to a collision. Figure 3(a) shows the resulting path of the two robots while Fig. 3(b) shows the distance of the two CoMs over time. As can be seen, both robots deviate from their desired path to avoid collision, showing a coordinated behavior. While being just a simple example this result highlights the potential of our MPC in solving collaborative tasks, further implementation in more complex scenarios, such as collaborative carrying is left for future work.
@@ IV-B Temporal Parallelization
In this subsection, we show the benefit of our implementation against the state-of-the-art solvers acados [8] and crocoddyl [5], in terms of average solving time for increasing horizon length. For this assessment, we initialized the robot in a random position and joint configuration and tasked the controller to perform a forward trot at a speed of $0.5$m/s. Acados uses the same cost and dynamics formulation as in our controller, while for the comparison with crocoddyl we adapted our reference generator and cost to match the reference used in the example provided with crocoddyl’s libraries. As reported in Fig. 4(a) and 4(b), our implementation shows faster update rates than both solvers in almost all tested scenarios. For the assesment against acados, with the SRBD model, our implementation is always faster. For the case of the WB formulation, crocoddyl is able to outperform our formulation with a prediction horizon smaller than $80$ nodes. This last outcome is primarily attributed to the overhead involved in transferring data to and from the GPU, which, for smaller problems, can be detrimental.
Noticeably, our formulation can achieve an update rate of 50Hz for a horizon length of 300 nodes.
@@ IV-C State Parallelization
Figure 5: Our implementation is able to control multiple robots in the same optimization loop, a typical scenario found in centralized controllers. In blue, we depict the solution time of acados, while in solid orange, ours, both considering the SRBD model. Furthermore, in dashed orange, we report the result obtained by our implementation using the WB model.
@@
—- $n_{\text{env}}$ Control Frequency Real-Time Factor
SRBD-MPC 4096 50Hz 370X
SRBD-MPC 4096 25Hz 570X
WB-MPC 4096 50Hz  35X
WB-MPC 4096 25Hz  75X
TABLE I: Performance metrics
@@ IV-D Model-Based Learning
Figure 6: Comparison of the solving time achieved by our implementation against the batch-parallelized solver feature of acados, when varying the number of environments being simulated in parallel. In this example, each environment has one robot that is controlled by a dedicated MPC.
In this subsection, we highlight the vectorization capabilities of our controller. Thanks to its full implementation in JAX, we can easily vectorize the computation of the MPC. Such ability is critical in the training scenario as shown in [26] or at runtime as shown in [40]. In Fig. 6, we compare our approach to the batched version of acados, which uses OpenMP to parallelize the computation of independent MPC instances on the CPU. As the plot shows, acados is competitive with the proposed MPC only for small batch sizes. This limitation comes from the reduced number of cores available on a CPU compared to the parallelization capabilities of modern GPUs.
@@ V Conclusion
This work presents a novel GPU-accelerated MPC framework implemented for legged robot locomotion. By leveraging the parallel processing capabilities of GPUs and implementing a Primal-Dual iLQR solver in JAX, we achieve logarithmic scaling in horizon length and linear scaling with state and control dimensions. Our approach demonstrates significant improvements over state-of-the-art solvers, achieving higher computational efficiency and scalability. This allows for the optimization of centralized controllers for multiple robots and the integration of large-scale parallel environments, enhancing learning-based control frameworks. Future works will include the exploitation of the presented MPC as a bias in the learning process of a locomotion policy. Furthermore, we would like to include a more sophisticated methodology for dealing with inequality constraints.

### v2

@@ III Model Predictive Control
In this section, we describe briefly the dynamics model $\boldsymbol{h}(\boldsymbol{x}_{i},\boldsymbol{u}_{i})$ presented in (1b), introducing two different models used for the formulation of the OCP presented in the result Sec. IV. The efficacy of the control input obtained from a MPC strategy is directly related to the rate at which such a control strategy can be re-planned and to the accuracy of its predictions, as shown by [solution_accuracy]. Therefore, we consider two main formulations from the literature: one based on the Single Rigid Body Dynamics (SRBD) model and one based on the Whole Body (WB) model. The first approximates the system as a single body with constant inertia, discarding the limb joint positions and the inertial variation from different limb configurations. This model is particularly effective in quadruped robots, where the limb mass is, in most cases, designed to be less than 10% of the total mass [cheeta_design]. On the other hand, the latter completely captures the dynamics of the robot, enabling the maximum exploitation of the system’s capabilities. This benefit comes at the cost of handling a larger state space and highly non-convex models, thus affecting the maximum frequency at which the optimization problem can be solved. Moreover, the SRBD-MPC state is limited to the center of mass linear and angular positions and velocities. This means that, when representing the robot’s orientation using quaternions, the state lies in $\mathbb{R}^{7}$. For such model, the control input consists only of the Ground Reaction Forces (GRF), and thus lies in $\mathbb{R}^{3n_{c}}$, where $n_{c}$ is the number of contact points, meaning that, for the evaluations with the quadruped, in Sec. IV $n$ is in $\mathbb{R}^{13}$ and $m$ in $\mathbb{R}^{12}$.
In contrast, the WB-MPC also includes joint positions and velocities in the state, and joint torques in the control input. This increases the state dimensionality to $\mathbb{R}^{13+2n_{\text{joint}}}$ and the control input dimensionality to $\mathbb{R}^{3n_{c}+n_{\text{joint}}}$, that translate for the quadruped model in Sec. IV in $n$ in $\mathbb{R}^{37}$ and $m$ in $\mathbb{R}^{24}$.
@@ IV Result
Figure 2: Comparison of the MPC optimal cost, evaluated from the same initial condition with acados [acados] (in blue) and with our implementation (in orange). Both solvers are evaluated at each call of the MPC and the costs are recorded.
Figure 3: Path of the two robots tracking their desired trajectories that will lead to a collision. We visualize the distance between the two robots with dashed gray lines of different shades, the darker the smaller distance.
Figure 4: Breakdown of the solving times for our implementation, acados (SQP) and crocoddyl (FDDP).
In this section, we present the results of various analyses that highlight the strength of the proposed approach. We evaluate the performance of our algorithm using the two models introduced in Sec. III (the SRBD and WB models). To benchmark our implementation against state-of-the-art solvers, we used acados [acados] (SQP) for comparisons on the SRBD model and crocoddyl [crocoddyl] (DDP) for the WB model. We used acados for the SRBD comparison, as it could be easily adapted for the evaluations with multiple robots, shown in Fig. 6 and 7. On the other hand, we used crocoddyl for the horizon length comparison with the WB model, as it showed significantly better performance than acados with the more complex model as shown in Fig. 4.
Our analysis includes investigating how the average solving time scales with the horizon length across all models and solvers. Additionally, we assess the solver’s performance as the system state dimension increases, as well as the vectorization capabilities of our approach. All the solving time comparisons consider a single iteration for each algorithm. The tests were conducted on a desktop computer equipped with an Intel Core i7-13700KF and an NVIDIA RTX 3080. We used Mujoco XLA, to evaluate the WB dynamics. In all the presented benchmarks, if not stated differently, we used the 15 kg Unitree Go2 robot, a torque-controlled quadruped platform, with a horizon length N of 50 nodes. As Fig. 1 and the accompanying video show, our WB-MPC can also be implemented for real-time control of other robot morphologies, like humanoids.
Figure 5: Average solving times when varying the horizon length (number of nodes) for the SRBD model (top) and WB model (bottom). In blue and green, we plot the solution time of two different state-of-the-art solvers (respectively acados [acados] and crocoddyl [crocoddyl]), while in orange, we plot the one of our implementation.
@@ IV-A Performance evaluation
First, we evaluate the ability of our solver to generate complex dynamic maneuvers. In this scenario, we are generating a barrel roll motion with the Aliengo robot, as shown in Fig. 1. Thanks to its multiple shooting formulation, we can initialize the solver with an infeasible initial guess, and the solver converges in just $27$ iterations to a solution that completes the task. We also assess the performance of the proposed solver in terms of solution optimality. In Fig. 2, we compared the values of the cost from the optimal solution given by our implementation against the one given by acados, which is chosen for this comparison because it allows using the same cost function as the one in our formulation. In this scenario, we are using the whole-body model and are controlling the robot in simulation at 50Hz. The robot is tasked to trot with a forward speed of $0.3$ m/s while being randomly pushed by an external disturbance of 5050N for $0.25$s. At every control loop, we record the value of the cost functional from both solvers, comparing solutions that start from the same initial condition. Figure 2 reports the values recorded during a $2.5$s time span. Although the two solutions are comparably similar, our implementation results in a $20\%$ reduction in the average cost along the trajectory.
To finally validate the MPC, we also performed experiments on the real hardware, as shown in Fig. 1 and in the accompanying video. For the real experiments, the solver runs on an Nvidia RTX 4050 laptop GPU. Figure 1 shows snapshots of the robot walking blindly up the steps and through uneven terrain using only proprioceptive data for state estimation [muse]. We also evaluated the proposed approach when used to control two different robots at the same time in a centralized control implementation in simulation. In this example, the two robots are controlled by the same MPC that evaluates the control input for both systems at 5050Hz. The MPC also includes a collision avoidance constraint, in the form of a quadratic penalty term. In the tested scenario, the two robots are tasked to track two perpendicular trajectories that cross each other and would lead to a collision. Figure 3 shows the resulting path of the two robots, where the dashed grey lines represent the distance between the two robots. As shown, both robots deviate from their desired paths to avoid collisions, demonstrating coordinated behavior. While being just a simple example, this result highlights the potential of our MPC in solving collaborative tasks, further implementation in more complex scenarios, such as collaborative carrying is left for future work. Figure 4 presents the average solving time breakdown for the three solvers in the quadruped trotting scenario with a horizon of N=100. It highlights the sequential backward pass in crocoddyl, the QP solver call in acados, and the parallel scan in our method for evaluating the search direction. A minor computational overhead of approximately $1ms$ arises from GPU communication, as only the initial state is sent and the commanded torques, joint positions, and velocities are retrieved.
@@ IV-B Temporal Parallelization
In this subsection, we show the benefit of our implementation against the state-of-the-art solvers acados [acados] and crocoddyl [crocoddyl], in terms of average solving time for increasing horizon length. For this assessment, we initialized the robot in a feasible random pose and joint configuration and tasked the controller to perform a forward trot at a speed of 0.50.5m/s. Acados uses the same cost and dynamics formulation as in our controller, while for the comparison with crocoddyl we adapted our reference generator and cost to match the reference used in the example provided with crocoddyl’s libraries. As reported in Fig. 5, our implementation shows faster update rates than both solvers in almost all tested scenarios. For the assesment against acados, with the SRBD model, our implementation is always faster. For the case of the WB formulation, crocoddyl is able to outperform our formulation with a prediction horizon smaller than $80$ nodes. This last outcome is primarily attributed to overhead involved in transferring data to and from the GPU, which, for smaller problems, can be detrimental. Noticeably, our formulation can achieve an update rate of 50Hz for a horizon length of 200 nodes.
@@ IV-C State Parallelization
Figure 6: Our implementation can control multiple robots in the same optimization, a typical scenario found in centralized controllers. In green, we depict the solution time of acados, while in orange, ours, considering the SRBD model. Furthermore, in dashed orange, we report the result obtained by our implementation using the WB model.
@@
Figure 7: Comparison of the solving time achieved by our implementation against the batch-parallelized solver feature of acados, when varying the number of environments being simulated in parallel. In this example, each environment has a dedicated MPC.
$n_{\text{env}}$ Control Frequency Real-Time Factor
[…]
TABLE I: Performance metric
@@ IV-D Model-Based Learning
In this subsection, we highlight the vectorization capabilities of our controller. Thanks to its full implementation in JAX, we can easily vectorize the computation of the MPC. Such ability is critical in the training scenario as shown in [DTC] or at runtime as shown in [mcts]. In Fig. 7, we compare our approach to the batched version of acados, which uses OpenMP to parallelize the computation of independent MPC instances on the CPU. As the plot shows, acados is competitive with the proposed MPC only for small batch sizes. This limitation comes from the reduced number of cores available on a CPU compared to the parallelization capabilities of modern GPUs.
@@ V Conclusion
This work presents a novel GPU-accelerated MPC framework implemented for legged robot locomotion. By leveraging the parallel processing capabilities of GPUs and implementing a Primal-Dual iLQR solver in JAX, we achieve logarithmic scaling in horizon length and square-log scaling with state and control dimensions. Our approach demonstrates significant improvements over state-of-the-art solvers, achieving higher computational efficiency and scalability. This allows for the optimization of centralized controllers for multiple robots and the integration of large-scale parallel environments, enhancing learning-based control frameworks. Future works will include the exploitation of the presented MPC as a bias in the learning process of a locomotion policy. Furthermore, we would like to include a more sophisticated methodology for dealing with inequality constraints.